OpenStudy (anonymous):

square root of 5b^2/125

OpenStudy (anonymous):

\[\sqrt{5b^2/125} \] first take out the b^2 square since the square root is b now you're left with \[b \sqrt{5/125}\] here we find two numbers that multiply together give you 125 but one has to have a square root what are they? 25*5 whats the square root of 25?! 5! so now \[5b \sqrt{5/5}\] whats 5/5?... 1? so what is the square root of 1? What do you have as a result?

OpenStudy (anonymous):

thank you b/5

OpenStudy (anonymous):

nope 5b remember \[5b \sqrt{5/5}\]